# Impact of dark states on the stationary properties of quantum particles with off-centered interactions in one dimension

### v1

@@ I Introduction
For single-component bosons, the $N$-body Hamiltonian with two-boy contact potential
@@
has been exhaustively studied for various trapping potentials $V(x)$. In the case of an infinite square well or no trapping potential, the model (1) is the Lieb-Liniger model Lieb and Liniger (1963) and solvable in the bosonic sector by the Bethe-ansatz for any interaction strength $g$ Gaudin (2014). Similarly, for $N=2$ and a harmonic trapping potential the model is also solvable for any $g$ Busch et al. (1998). The model (1) can also be generalized to multicomponent boson and fermion models Sowiński and Ángel García-March (2019); García-March et al. (2014); Harshman (2016a, b), for which the previous special cases are also solvable Sutherland (2004).
@@
The trap introduces another length scale to the problem, the harmonic oscillator length $a_{\rm{ho}}=\sqrt{\hbar/(m\omega)}$. We show below how the competition between the length scale $c$ and $a_{\rm{ho}}$ separates the energy spectrum into three regimes: (1) the exclusion regime for $c<a_{\rm{ho}}$ and low energies, where the model behaves similar to the contact interaction model, including two-body exclusion of particles for large $g$; (2) the truncation regime for $c>a_{\rm{ho}}$ and low energy, where the interaction length scale $c$ is so large that it suppresses the tail of the wave functions and creates a bunching effect; and (3) the crossover regime for $c\sim a_{\rm{ho}}$ where exclusion and bunching compete, and small variations of $c$ can lead to dramatic changes in the wave function variance. The interfaces between these regions are defined by the appearance of dark states to the interaction $U(g,c)$. Unlike the contact interaction, these dark states are both fermionic and bosonic and are sporadically distributed throughout the spectrum for specific values of $c$ corresponding to zeros of Hermite polynomials. At these points triple degeneracy occurs at infinite interactions, where two bosonic (fermionic) states cluster with another fermionic (bosonic) one. Moreover, dark states determine the competition between bunching and exclusion in the crossover regime.
The outline is as follows: in Sect. II, we analyze the symmetries of the model and identify the bosonic and fermionic sectors. In Sect. III, we provide analytic solutions for arbitrary $g$ and $c$ and analyze how the spectrum varies with the parameters. In Sect. IV, we construct the solutions at infinite interactions, analyze the corresponding energy spectrum, and subsequently discuss the ensuing symmetries. Finally, in Sect. V we summarize our results and examine possible realizations of the presented model.
II Structure and Symmetries of the Model
@@
The model (3) therefore is equivalent to two one-dimensional quantum systems. It is trivially integrable, with the role of the integral of motion for each degree of freedom played by each sub-Hamiltonian energy. Another consequence is that, like all 1D quantum systems with non-singular potentials, the spectrum of each sub-Hamiltonian should be singly-degenerate except in the case $g\to\infty$ Loudon (1959); Harshman (2017a). In what follows, we employ harmonic oscillator units such that $\hbar$, $m$, and $\omega$ all equal $1$.
@@
with energy $(n+1/2)$. As we show in the next section, solutions for arbitrary values of $g$ and $c$ can be found analytically by solving the transcendental equation given by matching boundary conditions for parabolic cylindrical functions at $x=\pm c$.
@@
Defining $\hat{\Pi}$ and $\hat{\Sigma}$ as operators that represent the generators $\Pi$ and $\Sigma$ acting on the Hilbert space of wave functions, the states $\Phi_{N}(X)$ and $\phi_{n}(x)$ transform as:
@@
The center-of-mass wave functions are invariant under particle exchange, but the relative wave functions can be split into two sectors: bosonic wave functions that are even under exchange and fermionic wave functions that are odd under exchange. For the case of two particles, we also see that particle exchange eigenstates coincide with relative parity eigenstates: bosonic wave functions are necessarily even under (relative) parity and fermionic are odd under relative parity. These sectors decouple and can be considered separately.
@@
The symmetries in the case $g\to\infty$ will be discussed below in Sect. IV.
III Generic Interaction Strength
@@
In this section, we provide expressions for the solutions $\phi_{n}(x)$ of the relative sub-Hamiltonian $H_{\mathrm{rel}}(g,c)$ when the interaction strength $g$ is finite. We will also identify two critical features of the model: $(i)$ the role played by dark states, i.e., states of the non-interacting Hamiltonian whose wave function nodes coincide with the interaction displacement $c$ and thus remain unperturbed; and $(ii)$ the role played by the competition between the harmonic oscillator length scale and the interaction offset scale $c$.
III.1 Solutions of the Relative Sub-Hamiltonian
@@
The first step towards finding the eigenspectrum of the relative sub-Hamiltonian Eq. (II) for generic $g$ is to divide the relative $x$ coordinate into three regions $I=(-\infty,-c)$, $I\!I=(-c,c)$ and $I\!I\!I=(c,\infty)$. The solutions in these intervals are known: they are linear combinations of parabolic cylinder functions $D_{Q}(x)$ Abramowitz and Stegun (1948); Aouadi et al. (2016); Avakian et al. (1987),
@@
where $Q_{n}\equiv\epsilon_{n}-1/2$ and the coefficients are yet to be determined. In the case where $Q_{n}=n$, an integer, these functions reduce to the $n$-order Hermite polynomials Abramowitz and Stegun (1948),
$D_{n}(x)=2^{-n/2}H_{n}(x/\sqrt{2})e^{-1/4x^{2}}=(2\pi)^{1/4}\sqrt{n!}\,\phi^{0}_{n}(x).$ (12)
Note that since we are interested in bound state solutions due to the harmonic trap, we demand that the eigenstates vanish at $x\to\pm\infty$. From the asymptotic expansions Abramowitz and Stegun (1948); Aouadi et al. (2016) $D_{Q_{n}}(x\to\infty)\sim e^{-x^{2}/4}x^{Q_{n}}$ and $D_{Q_{n}}(x\to-\infty)\sim\sqrt{2\pi}/\Gamma(-Q_{n})e^{x^{2}/4}x^{-1-Q_{n}}$ we infer that $\alpha_{I}=\beta_{I\!I\!I}=0$.
@@
Note that the above relation is general and the energy levels $\epsilon_{n}=Q_{n}+1/2$ are determined for arbitrary finite interaction strength $g$. The equation holds also for arbitrary displacement $c$ except from the special cases where $D_{Q_{n}}(\pm c)=0$, which are treated separately. These occur whenever $Q_{n}$ is an integer, and the zeros correspond to the roots of Hermite polynomials [Eq. (12)], and hence to non-interacting eigenstates $\phi^{0}_{n}(x)$. These states do not ‘feel’ the interaction, i.e. they are dark states. This behavior is similar to that of fermionic states, which do not ‘feel’ a contact interaction; see below. We will see evidence of these dark states in the spectrum below.
@@
III.2 Competition between scales
Now we analyze how the competition between the oscillator length scale ($1$ in our units) and the interaction displacement $c$ is imprinted on the energy levels and eigenstates. The dark states described above also leave evidence in the spectrum.
First, consider the displacement $c=0.75<1$, smaller than the trap length scale. The corresponding energy level structure is presented in Fig. 1. As expected due to the non-zero displacement, fermionic energy levels are also affected by the interactions [dashed lines in Fig. 1(a)], contrary to the case of zero-range interactions Busch et al. (1998); Budewig et al. (2019). There is however one energy level ($n=4$) which barely depends on the interaction strength, i.e., a dark state. This occurs because the first root of the corresponding Hermite polynomial $H_{4}(x/\sqrt{2})$ lies very close to the displacement $c$  Salzer et al. (1952). As $g\to\infty$, the energy levels of the adjacent bosonic ($n=2$) and fermionic ($n=3$) state tend asymptotically to the energy of the dark state, signalling a triple degeneracy. The latter is present at attractive interactions as well, where the $n=4$ dark state now clusters with the $n=5,n=6$ levels.
Except from the triple degeneracy induced by the dark state, all other low energy levels are doubly degenerate as $g\to\infty$. Adjacent bosonic and fermionic energy levels cluster together in a way reminiscent of the Tonks-Girardeau regime, manifested in systems with strongly repulsive short-range interactions Tonks (1936); Girardeau (1960); Kehrberger et al. (2018). In that regime, bosons turn into hardcore particles avoiding each other and their energy levels become degenerate with those of non-interacting fermions. By analogy, we call this displacement parameter range ($c<1$) the exclusion regime.
The exclusion is demonstrated in the transition of the profiles of the first bosonic and fermionic eigenstates ($n=0,1$) from attractive to strongly repulsive interactions [Fig. 1(b), (c)]. The two particles tend to avoid the narrow region $I\!I$ and delocalize away from their interaction centers in regions $I$ and $I\!I\!I$. For strong repulsion ($g=20$), the probability to find the two particles inside the interval $(-c,c)$ is very small for both the bosonic and the fermionic state. Note that as the interactions become even more repulsive ($g=100$), the corresponding probability tends to zero.

### v3

@@ I Introduction
For single-component bosons, the $N$-body Hamiltonian with two-body contact potential
@@
has been exhaustively studied for various trapping potentials $V(x)$. In the case of an infinite square well or no trapping potential, the model (1) is the Lieb-Liniger model Lieb and Liniger (1963) and solvable in the bosonic sector by the Bethe-ansatz for any interaction strength $g$ Gaudin (2014); Yang (1967); Gaudin (1967). Similarly, for $N=2$ and a harmonic trapping potential the model is also solvable for any $g$ Busch et al. (1998). The model (1) can also be generalized to multicomponent boson and fermion models Sowiński and Ángel García-March (2019); García-March et al. (2014); Harshman (2016a, b), for which the previous special cases are also solvable Sutherland (2004).
@@
The trap introduces another length scale to the problem, the harmonic oscillator length $a_{\rm{ho}}=\sqrt{\hbar/(m\omega)}$. We show below how the competition between the length scale $c$ and $a_{\rm{ho}}$ separates the energy spectrum into three regimes: (1) the exclusion regime for $c<a_{\rm{ho}}$ and low energies, where the model behaves similar to the contact interaction model, including two-body exclusion of particles for large $g$; (2) the truncation regime for $c>a_{\rm{ho}}$ and low energy, where the interaction length scale $c$ is so large that it suppresses the tail of the wave functions and creates a bunching effect; and (3) the crossover regime for $c\sim a_{\rm{ho}}$ where exclusion and bunching compete, and small variations of $c$ can lead to dramatic changes in the wave function variance.
The interfaces between these regions are defined by the appearance of dark states to the interaction $U(g,c)$. These are non-interacting eigenstates that simultaneously solve the interacting problem Werner and Castin (2006); Werner (2008), and they are typically encountered in systems with contact interactions Busch et al. (1998). The particles residing in such states are insensitive to the tuning of the interaction strength, and we use the term dark states for them in analogy to states in quantum optics that are insensitive to optical driving Fleischhauer et al. (2005). Unlike the contact interaction, these dark states are both fermionic and bosonic and are sporadically distributed throughout the spectrum for specific values of $c$ corresponding to zeros of Hermite polynomials. At these points triple degeneracy occurs at infinite interactions, where two bosonic (fermionic) states cluster with another fermionic (bosonic) one. Moreover, dark states determine the competition between bunching and exclusion in the crossover regime.
The outline is as follows: in Sect. II, we analyze the symmetries of the model and identify the bosonic and fermionic sectors. In Sect. III, we construct the solutions at infinite interactions, analyze the corresponding energy spectrum, and subsequently discuss the ensuing symmetries. In Sect. IV, we provide analytic solutions for arbitrary $g$ and $c$ and analyze how the spectrum and probability density vary with the parameters. Finally, in Sect. V we summarize our results and examine possible realizations of the presented model.
II Structure of the Model
@@
The model (3) therefore is equivalent to two one-dimensional quantum systems. It is trivially integrable, with the role of the integral of motion for each degree of freedom played by each sub-Hamiltonian energy. Another consequence is that, like all 1D quantum systems with non-singular potentials, the spectrum of each sub-Hamiltonian should be non-degenerate except in the case $g\to\infty$ Loudon (1959); Harshman (2017a). In what follows, we employ harmonic oscillator units such that $\hbar$, $m$, and $\omega$ all equal $1$.
@@
with energy $(n+1/2)$. As we show in the next sections, solutions for arbitrary values of $g$ and $c$ can be found analytically by solving the transcendental equation given by matching boundary conditions for parabolic cylindrical functions at $x=\pm c$.
@@
Defining $\hat{\Pi}$ and $\hat{\Sigma}$ as operators acting on the Hilbert space of wave functions that represent the coordinate transformations $\Pi$ and $\Sigma$, the states $\Phi_{N}(X)$ and $\phi_{n}(x)$ transform as:
@@
The center-of-mass wave functions are invariant under particle exchange, but the relative wave functions can be split into two sectors: bosonic wave functions that are even under exchange and fermionic spatial wave functions that are odd under exchange. The two fermions are spin polarized, i.e., single component so that the combined spatial and spin relative wavefunctions are odd under the $\Sigma$ operation. For the case of two particles, we also see that particle exchange eigenstates coincide with relative parity eigenstates: bosonic wave functions are necessarily even under (relative) parity and fermionic are odd under relative parity. These sectors decouple and can be considered separately.
@@
The symmetries in the $g\to\infty$ case will be discussed after presenting the energy spectrum.
III Infinite interactions
Despite the fact that $g=\infty$ is a special regime requiring separate treatment, further insight can be gained regarding the structure of eigenstates and the degeneracy of energy levels. Such knowledge will be useful later on when analyzing the structure of eigenstates at finite interaction strengths.
III.1 Energy levels
To find the energy level structure at $g=+\infty$, we proceed as follows. The relative $x$ coordinate is separated into three regions, $I=(-\infty,-c)$, $I\!I=(-c,c)$ and $I\!I\!I=(c,\infty)$. These are disjoint intervals, given that there is effectively a hard wall at the interaction centers ±$\pm c$. Therefore, the wavefunctions have to vanish at the intersections of the intervals. The relative Hamiltonian in regions $I$ and $I\!I\!I$ corresponds to the Hamiltonian of a single particle confined in a harmonic oscillator, truncated by a hard-wall boundary at $x=-c$ ($I$) or $x=c$ ($I\!I\!I$). In region $I\!I$, $H_{\rm{rel}}$ corresponds to that of a single particle confined in a box potential superimposed with a harmonic oscillator.
We first focus on the latter region. The relative wave functions are known, they are combinations of parabolic cylinder functions Abramowitz and Stegun (1948); Aouadi et al. (2016); Avakian et al. (1987),
where $Q^{(in)}_{n}\equiv\epsilon^{(in)}_{n}-1/2$ and the coefficients are yet to be determined. The superscript denotes that the two particles are located inside the interval determined by the interaction centers, $(-c,c)$. In the case where $Q^{(in)}_{n}=n$, an integer, these functions reduce to the $n$-order Hermite polynomials Abramowitz and Stegun (1948),
$D_{n}(x)=2^{-n/2}H_{n}(x/\sqrt{2})e^{-x^{2}/4}=(2\pi)^{1/4}\sqrt{n!}\,\phi^{0}_{n}(x).$ (12)
The hard-wall boundary conditions imposed by the box potential imply that $\phi^{(in)}_{n}(-c)=\phi^{(in)}_{n}(c)=0$. These two equations can be cast in the form of a matrix eigenvalue problem with zero eigenvalue Consortini and Frieden (1976),
$D^{2}_{Q^{(in)}_{n}}(c)-D^{2}_{Q^{(in)}_{n}}(-c)=0.$ (14)
For the other two symmetric regions it suffices to solve the relative Hamiltonian in only one of them. This is due to the fact that $H_{\rm{rel}}$ is invariant under spatial inversion $\hat{\Pi}$, even for infinite interactions [see also Sec. II]. As a result, the energy levels in regions $I$ and $I\!I\!I$ coincide, denoted as $\epsilon^{(out)}_{\nu}$. Focusing on region $I\!I\!I$ for example, the wave functions read $\chi^{(out,I\!I\!I)}_{\nu}(x)=\alpha^{(out,I\!I\!I)}D_{Q^{(out)}_{\nu}}(x)$, where $Q^{(out)}_{\nu}\equiv\epsilon^{(out)}_{\nu}-1/2$. The superscript denotes that both particles are located outside of the interaction centers interval $(-c,c)$, and in particular in region $I\!I\!I$. Note that we are interested in bound state solutions, and thus only a single parabolic cylinder function is considered, since $D_{Q^{(out)}_{n}}(-x)$ diverges exponentially as $x\to\infty$ Abramowitz and Stegun (1948); Aouadi et al. (2016). Imposing the hard-wall boundary condition at $x=c$, the energy levels $\epsilon^{(out)}_{\nu}$ are determined Grosche (1993),
$D_{Q^{(out)}_{\nu}}(c)=0.$ (15)
The associated energy eigenstates $\chi^{(out,I\!I\!I)}_{\nu}(x)$ are not eigenstates of the transformations $\Pi$ and $\Sigma$ since they are mapped to interval $I$, e.g. $\hat{\Sigma}\chi^{(out,I\!I\!I)}_{\nu}(x)=\chi^{(out,I)}_{\nu}(x)$. Simultaneous eigenstates of the relative Hamiltonian and kinematic symmetries can be constructed:
From the above analysis it becomes clear that the bosonic and fermionic levels $\epsilon^{(out)}_{n}$ are doubly-degenerate, and that particles occupying these eigenstates are localized outside of the interval $(-c,c)$. On the other hand, the two particles are strictly found within $(-c,c)$ when the $\phi^{(in)}_{n}$ eigenstates are populated. In this case the bosonic and fermionic levels are distinct [Eq. (14)] and they are non-degenerate. Varying the displacement $c$ is equivalent to moving the hard walls, and thus shifting the energy levels $\epsilon^{(in)}_{n}$ and $\epsilon^{(out)}_{n}$ [Fig. 1]. These two kinds of eigenstates correspond to different Hamiltonians, and thus there are only exact crossings between them. The positions of these crossings correspond to roots of Hermite polynomials, where the non-interacting eigenstates vanish [Eq. (12)]. For these values of $c$, there are therefore ‘dark’ eigenstates of the relative Hamiltonian that have nodes exactly where the interaction lies. Therefore, these dark states indicate special points where there is a triple degeneracy.
The energy levels at $g=+\infty$ [Fig. 1] help us to neatly classify the density profiles of the two-particle system, depending on the value of $c$. In particular, $\epsilon^{(in)}_{0}$ and $\epsilon^{(out)}_{0,1}$ delineate three regimes. On the left of $\epsilon^{(in)}_{0}$, for $c\ll 1$, all eigenstates are doubly-degenerate, and particles are excluded from the interval $(-c,c)$. Adjacent bosonic and fermionic energy levels cluster together in a way reminiscent of the Tonks-Girardeau regime, manifested in systems with strongly repulsive short-range interactions Tonks (1936); Girardeau (1960); Kehrberger et al. (2018). In that regime, bosons turn into hardcore particles avoiding each other and their energy levels become degenerate with those of non-interacting fermions. By analogy, we call this displacement parameter range the exclusion regime.
On the right of $\epsilon^{(out)}_{0,1}$, $c\gg 1$, all eigenstates are non-degenerate. Particles are localized within the interaction center interval $(-c,c)$, and the energy levels saturate to their non-interacting values at large $c$. In this region, the interaction centers are located at the edges of the harmonic trap, and the two particles barely feel any interaction. The oscillator length is the only relevant length scale, dictating the exponential decay of the eigenstates at large separations. Since the energy spectrum resembles that of an harmonic oscillator, truncated at the edges due to the interaction centers, the regime of large interaction displacement is called the truncation region (T). In-between these two regions, non- and doubly-degenerate eigenstates coexist, denoting the crossover regime (C). When $c$ coincides with the roots of Hermite polynomials, triple degeneracy occurs, as $\epsilon^{(out)}_{n}$ match with $\epsilon^{(in)}_{n}$.
III.2 Symmetries
For finite $g$ and $c$, all energy levels of the relative sub-Hamiltonian are non-degenerate, as is true for any solutions to the 1D Schrödinger equation defined on a path-connected interval. More generally, any system whose maximal kinematic symmetry is Abelian should only be non-degenerate. How then can we understand the additional double and triple degeneracies that occur in the $g\to\infty$ limit?
These can be understood by recognizing that the domain of the relative coordinate $x$ becomes effectively disconnected for infinite $g$. The three domains, $I=(-\infty,-c)$, $I\!I=(-c,c)$, and $I\!I\!I=(c,\infty)$ act like an independent quantum system, each experiencing independent time evolution. Energy eigenstates are localized to each of these three distinct regions in the $g\to\infty$ limit. To see this, note that the relative sub-Hamiltonian $H_{\mathrm{rel}}(g,c)$ can be re-expressed as the direct sum of three Hamiltonians:
$\lim_{g\to\infty}H_{\mathrm{rel}}(g,c)=H_{I}\oplus H_{I\!I}\oplus H_{I\!I\!I},$ (18)
where $H_{R}$ is the 1D harmonic oscillator Hamiltonian restricted to region $R\in\{I,I\!I,I\!I\!I\}$. These restricted Hamiltonians commute, so instead of one time translation symmetry parameterized by $t\in T_{t}\sim\mathbb{R}$, there are now three parameterized by $(t_{I},t_{I\!I},t_{I\!I\!I})\in T_{t}^{\times 3}\sim\mathbb{R}^{3}$ and represented by products of the three unitary operators $\hat{U}_{R}(t_{R})=\exp(-iH_{R}t_{R})$. Equivalently, the phase difference between disjoint regions in the $g\to\infty$ limit is not an observable quantity Harshman (2017a). Only when $g$ is finite, is there coupling between adjacent regions that locks their relative phase.
Figure 1: (a) Energy levels at $g=+\infty$ versus the displacement $c$. The positions of the exact crossings between the doubly-degenerate $\epsilon^{(out)}_{n}$ and non-degenerate $\epsilon^{(in)}_{n}$ levels designate three regions, the exclusion (E), crossover (C), and truncation (T) region.
However, this additional symmetry of tripled time evolution is still an Abelian kinematic symmetry, and therefore not enough to explain the systematic double and triple degeneracies that occur for the $g\to\infty$ limit of $H_{\mathrm{rel}}(g,c)$. If regions $I$, $I\!I$, and $I\!I\!I$ were all intervals with inequivalent domains, then the spectrum would be non-degenerate except for so-called accidental degeneracies where two states (or even more rarely, three states) would coincide in energy. For our system, the regions $I$ and $I\!I\!I$ are equivalent intervals exchanged by relative parity $\Sigma$. Therefore the spectrum of $H_{I}$ and $H_{I\!I\!I}$ coincide and the out states $\chi_{\nu}^{(out,I)}$ and $\chi_{\nu}^{(out,I\!I\!I)}$ are double-degenerate pairs that can be symmetrized into bosonic or antisymmetrized into fermionic states $\phi_{n}^{(out)}$.
$\hat{\Sigma}\hat{U}_{I}(t_{I})=\hat{U}_{I\!I\!I}(t_{I\!I\!I})\hat{\Sigma}.$ (19)
Further, triple degeneracies occur when states in the spectrum of region $I\!I$ have energies that coincide with states of the outer regions $I$ and $I\!I\!I$. These spectral ‘accidents’ occur precisely when the displacements ±$\pm c$ align with the $n$th Hermite polynomial root, $H_{n}(c/\sqrt{2})=0$. When $n$ is even, the state in region $I\!I$ is even, and the triply-degenerate state is comprised of two bosonic states (one inside and one outside the central region) and one fermionic state. When $n$ is odd, triple degeneracy occurs between two fermionic states and one bosonic. Understanding these accidental degeneracies as the result of dark states provides an interesting perspective on an old subject Jauch and Hill (1940); McIntosh (1959); Louck and Metropolis (1981); Moshinsky and Quesne (1983); Leyvraz et al. (1997), but not all accidental degeneracies can be understood this way.
IV Generic Interaction Strength
@@
In this section, we provide expressions for the solutions $\phi_{n}(x)$ of the relative sub-Hamiltonian $H_{\mathrm{rel}}(g,c)$ when the interaction strength $g$ is finite. We will also identify two critical features of the model: $(i)$ the role played by dark states, i.e., states of the non-interacting Hamiltonian whose wave function nodes coincide with the interaction displacement $c$ and thus remain unperturbed; and $(ii)$ the transitions occurring around the displacement $c$ where triple degeneracy manifests at $g=\infty$.
IV.1 Solutions of the Relative Sub-Hamiltonian
@@
The first step towards finding the eigenspectrum of the relative sub-Hamiltonian Eq. (II) for generic $g$ is to divide the relative $x$ coordinate into three regions $I=(-\infty,-c)$, $I\!I=(-c,c)$ and $I\!I\!I=(c,\infty)$, similarly to the case where $g=\infty$. The solutions in these intervals are linear combinations of parabolic cylinder functions $D_{Q}(x)$ Abramowitz and Stegun (1948); Aouadi et al. (2016); Avakian et al. (1987),
@@
where $Q_{n}\equiv\epsilon_{n}-1/2$ and the coefficients are yet to be determined.
Since we are interested in bound state solutions due to the harmonic trap, we demand that the eigenstates vanish at $x\to\pm\infty$. From the asymptotic expansions Abramowitz and Stegun (1948); Aouadi et al. (2016) $D_{Q_{n}}(x\to\infty)\sim e^{-x^{2}/4}x^{Q_{n}}$ and $D_{Q_{n}}(x\to-\infty)\sim\sqrt{2\pi}/\Gamma(-Q_{n})e^{x^{2}/4}x^{-1-Q_{n}}$ we infer that $\alpha_{I}=\beta_{I\!I\!I}=0$.
@@
Note that the above relation is general and the energy levels $\epsilon_{n}=Q_{n}+1/2$ are determined for arbitrary finite interaction strength $g$. The equation holds also for arbitrary displacement $c$ except from the special cases where $D_{Q_{n}}(\pm c)=0$, which are treated separately. These occur whenever $Q_{n}$ is an integer, and the zeros correspond to the roots of Hermite polynomials [Eq. (12)], and hence to non-interacting eigenstates $\phi^{0}_{n}(x)$. These states do not ‘feel’ the interaction, i.e. they are dark states. This behavior is similar to that of fermionic states, which do not ‘feel’ a zero-range contact interaction; see below. We will see evidence of these dark states in the spectrum below.
Note that irrespective of the interaction strength, there always exist high-lying excited states which are barely affected by the $\delta$ potentials, since the interaction energy is small. In our analysis, however, these are not taken into account, and the dark states always refer to low-lying energy eigenstates.
@@
IV.2 Competition between scales
Now we analyze how a finite interaction strength affects the energy spectrum and the resulting eigenstates, especially close to the triple degeneracy points identified at $g=\infty$.
First, consider the displacement $c=0.75<1$, inside the exclusion regime [Fig. 1]. The corresponding energy level structure is presented in Fig. 2. As expected due to the non-zero displacement, fermionic energy levels are also affected by the interactions [dashed lines in Fig. 2(a)], contrary to the case of zero-range interactions Busch et al. (1998); Budewig et al. (2019). There is however one energy level ($n=4$) which barely depends on the interaction strength, i.e., a dark state. This occurs because the first root of the corresponding Hermite polynomial $H_{4}(x/\sqrt{2})$ lies very close to the displacement $c$  Salzer et al. (1952). As $g\to\infty$, the energy levels of the adjacent bosonic ($n=2$) and fermionic ($n=3$) state tend asymptotically to the energy of the dark state, signalling a triple degeneracy, as seen also from Fig. 1. The latter is present at attractive interactions as well, where the $n=4$ dark state now clusters with the $n=5,n=6$ levels. Except from the triple degeneracy induced by the dark state, all other low energy levels are doubly degenerate as $g\to\infty$, given that this is the exclusion regime.
The exclusion from the $(-c,c)$ interval is demonstrated in the transition of the profiles of the first bosonic and fermionic eigenstates ($n=0,1$) from attractive to strongly repulsive interactions [Fig. 2(b), (c)]. The two particles tend to avoid the narrow region $I\!I$ and delocalize away from their interaction centers in regions $I$ and $I\!I\!I$. For strong repulsion ($g=20$), the probability to find the two particles inside the interval $(-c,c)$ is very small for both the bosonic and the fermionic state. Note that as the interactions become even more repulsive ($g=100$), the corresponding probability tends to zero. This behavior occurs because the wavefunctions start to resemble the $\phi^{(out)}_{n}(x)$ eigenstates occurring at $g=\infty$, which are supported only outside of the $(-c,c)$ interval.